# A homogeneous geometry of low-rank tensors

### v1

@@ Abstract
We consider sets of fixed CP, multilinear, and $T$ rank tensors, and derive conditions for when (the smooth parts of) these sets are smooth homogeneous manifolds. For CP and $T$ ranks, the conditions are essentially that the rank is sufficiently low. These homogeneous structures are then used to derive Riemannian metrics whose geodesics are both complete and efficient to compute.
@@ 1 Introduction
Consider the group action that multiplies tensors in each mode by an invertible matrix, which can also be seen as a change of basis in each mode. While there are several ways to define the rank of a tensor, for almost all of them, the rank is invariant under this action. There are in general two obstacles to using it to identify sets of fixed-rank tensors as homogeneous manifolds:
@@
In this paper, we consider three notions of rank: canonical polyadic (CP) rank, tensor train ($T$) rank, and multilinear rank. The two obstacles are overcome by:
@@
We thus identify three families of fixed-rank tensor manifolds as smooth homogeneous manifolds. For each of these, there is a so-called canonical Riemannian metric. We show how the geodesics in this metric can be computed efficiently.
@@
In low-rank approximation, the manifold perspective has been used to study several different fixed-rank tensor spaces. We mention here CP rank [Breiding18], $T$ rank [Holtz12, Uschmajew20], hierarchical Tucker rank [Uschmajew13], and multilinear rank [Koch10].
In [Uschmajew20], it is mentioned that “the set of all tensors with canonical rank bounded by $k$ is typically not closed. Moreover, while the closure of this set is an algebraic variety, its smooth part is in general not equal to the set of tensors of fixed rank $k$ and does not admit an easy explicit description.”. Our contribution is to find an “easy explicit description” for a subset of those manifolds.
We also mention that fixed-rank tensor manifolds have, to our knowledge, not previously been equipped with a Riemannian metric with known geodesics, other than in the rank 11 case [Swijsen22b, Jacobsson24].
In Riemannian optimization, line searches and gradient descents are often performed along geodesics, or along approximate geodesics, called retractions. As already mentioned, on any homogeneous manifold there is a distinguished Riemannian metric called the canonical metric. This metric has two big advantages: geodesics on a quotient can be described in terms of geodesics on its numerator; and those geodesics are always complete, meaning they can be extended to arbitrary length. Similarly, a retraction on the quotient can be induced by a retraction on the numerator. Riemannian optimization using the canonical metric has been explored, for example, on the Grassmann [Helmke07, Sato14, Boumal15, Bendokat24], Stiefel [Edelman98, Li20, Gao22], and symplectic Stiefel [Gao21, Bendokat21] manifolds.
@@
Section˜2 briefly recaps the concepts from three different fields of study—homogeneous spaces, multilinear algebra, and algebraic geometry—that we use to construct our homogeneous structure. Sections˜3, 4, and 5 apply these to sets of tensors with fixed CP rank, multilinear rank, and $T$ rank respectively. The constructions are similar to each other, and the basic results are repeated without proof in sections˜4 and 5 after having been introduced in detail in section˜3.
@@ 2.3 Multilinear algebra
The CP rank (often just rank) of a tensor $T\in\mathbb{R}^{n_{1}}\otimes\dots\otimes\mathbb{R}^{n_{d}}$ is the smallest $r$ such that $T$ is a sum of $r$ rank 11 tensors:
@@
Figure 2: Penrose diagram for the CP decomposition. Edges are labeled with the vector space they represent, and $V_{i}=\begin{bmatrix}v_{i}^{1}&\dots&v_{i}^{r}\end{bmatrix}$. II is the diagonal tensor $I_{i_{1}\dots i_{d}}=1$ if $i_{1}=\dots=i_{d}$ else $I_{i_{1}\dots i_{d}}=0$.
@@
The $T$ rank of a tensor $T\in\mathbb{R}^{n_{1}}\otimes\dots\otimes\mathbb{R}^{n_{d}}$ is the tuple $(s_{1},\dots,s_{d-1})$ with
@@
If $T\in\mathbb{R}^{n_{1}}\otimes\dots\otimes\mathbb{R}^{n_{d}}$ has $T$ rank $(s_{1},\dots,s_{d-1})$ then $T$ can be written as
@@
Such an expression is called a $T$ decomposition. It is illustrated in figure˜4.
Figure 4: Penrose diagram for the $T$ decomposition ˜2.10. Edges are labeled with the vector space they represent.
Note also that a converse of proposition˜5 is true, that a $T$ decomposition with inner dimensions $s_{1}$, …, $s_{d-1}$ has $T$ rank at most $(s_{1},\dots,s_{d-1})$.
@@ 3 The CP manifold
Note that $v_{i}^{1}$, …, $v_{i}^{r}$ are linearly independent in $\mathbb{R}^{n_{i}}$ implies $r\leq n_{i}$ for all ii.
@@ 3.1 Smooth manifold
We now consider the subgroup of $G$ that fixes the $T$ defined in ˜$3.$ Applying (Kruskal’s theorem)., we have the following result.
@@ 3.2 Representatives
and note that $h=(h_{1},\dots,h_{d})\in H$. Here, we see that $g_{11}$ needs to be invertible. If $g_{11}$ is not invertible, there is a permutation matrix $P$ such that $g_{i}^{\prime}=Pg_{i}$ has an invertible block $g^{\prime}_{11}$. In this way we only need to specify $g_{11}$, $g_{21}$, and $P$ for each ii.
@@ 3.3 Riemannian manifold
For any $g\in G$, the coset gHgH is a submanifold of $G$. Its tangent space at $g$ is called the vertical space, and is denoted $\mathcal{V}_{g}$. The orthogonal complement to the vertical space is called horizontal space, and is denoted $\mathcal{H}_{g}$. Thus $\mathfrak{h}$ and $\mathfrak{m}$ are the vertical and horizontal spaces at 11.
@@
We now want to derive a more explicit description of the horizontal space. First, note that $\mathcal{V}_{g}=g\mathfrak{h}$, so that the vertical space consists of elements $Y=(Y_{1},\dots,Y_{d})$ where the $Y_{i}=g_{i}\begin{bmatrix}Y_{11}&Y_{12}\\ &Y_{22}\end{bmatrix}$. From now on, we suppress the $i$ superscript, but note that $Y_{11}=Y^{i}_{11}$ and the other blocks still depends on ii. Second, if we choose a block lower triangular representative as in section˜3.2, we have
@@
But BABA is an $r\times r$ matrix, allowing 1/(1+BA)-1/(1+BA) to be evaluated efficiently. This trick is also useful later for evaluating the matrix exponential. For $\Gamma_{12}$, we have the following expression,
@@
has the same dot product with the $k$th column of $\begin{bmatrix}g_{11}\\ g_{12}\end{bmatrix}$ for every ii.
@@ 3.4 Riemannian homogeneous manifold
Conversely, assume there exists an invariant subspace $\mathfrak{p}$ when $n_{i}>r$ for some ii. Note that elements in $\mathfrak{p}$ are determined by their projection to $\mathfrak{m}$. Concretely, elements in $\mathfrak{p}$ are determined by their $r$ first columns. So consider an element $X=(X_{1},\dots,X_{d})\in\mathfrak{p}$ with
@@ 4.1 Smooth manifold
Let $E_{i}$ be the identity matrix $I_{n_{i}\times t_{i}}$ seen as an element of $\mathbb{R}^{n_{i}}\otimes\mathbb{R}^{t_{i}}$ and let II be the identity matrix $I_{t_{1}\times t_{1}}$ seen as an element of $\mathbb{R}^{t_{1}}\otimes\dots\otimes\mathbb{R}^{t_{d}}$. Then define
@@ 5 The tensor train manifold
The $T$ decomposition is unique up to a change of basis in $\mathbb{R}^{s_{1}}$, …, $\mathbb{R}^{s_{d-1}}$. More precisely,
@@
Uschmajew and Vandereycken [Uschmajew13, proposition 3] shows that this holds for Hierarchical Tucker decompositions, of which the $T$ decomposition is a special case.
@@ 5.1 Smooth manifold
Let $\operatorname{ttrank}^{-1}(s_{1},\dots,s_{d-1})$ denote the set of tensors with $T$ rank $(s_{1},\dots,s_{d-1})$. Like $\operatorname{rank}^{-1}(r)$, its closure is an algebraic variety.
@@
From ˜2.9, it is clear that $T$ rank is preserved under $G$. The action is thus well-defined.
@@
The set of tensors with $T$ rank $(s_{1},\dots,s_{d-1})$ and multilinear rank $(s_{1},$ $s_{1}s_{2},\dots,s_{d-1})$ is a smooth homogeneous manifold,
@@ Appendix B Appendix: Counting the operations
  • no extra operations to build AA,
@@
    • $2n_{i}r^{2}$ operations to multiply with AA from the left,
@@
    • $2n_{i}r^{2}$ operations to multiply with AA from the left.

### v2

@@ Abstract
We consider sets of fixed CP, multilinear, and TT rank tensors, and derive conditions for when (the smooth parts of) these sets are smooth homogeneous manifolds. For CP and TT ranks, the conditions are essentially that the rank is sufficiently low. These homogeneous structures are then used to derive Riemannian metrics whose geodesics are both complete and efficient to compute.
@@ 1 Introduction
Consider the group action that multiplies tensors in each mode by an invertible matrix, which can also be seen as a change of basis in each mode. While there are several ways to define the rank of a tensor, for most of them, the rank is invariant under this action. There are in general two obstacles to using it to identify sets of fixed-rank tensors as homogeneous manifolds:
@@
In this paper, we consider three notions of rank: canonical polyadic (CP) rank, multilinear rank, and tensor train (TT) rank. The two obstacles are overcome by:
@@
We are thus able to identify three families of fixed-rank tensor manifolds as smooth homogeneous manifolds. For each of these, there is a so-called canonical Riemannian metric. We show how the geodesics in this metric can be computed efficiently.
@@
In low-rank approximation, the manifold perspective has been used to study several different fixed-rank tensor spaces. We mention here CP rank [Breiding18], TT rank [Holtz12, Uschmajew20], hierarchical Tucker rank [Uschmajew13], and multilinear rank [Koch10].
In [Uschmajew20], it is mentioned that “the set of all tensors with canonical rank bounded by $k$ is typically not closed. Moreover, while the closure of this set is an algebraic variety, its smooth part is in general not equal to the set of tensors of fixed rank $k$ and does not admit an easy explicit description.”. We contribute an “easy explicit description” for a subset of those manifolds.
We also mention that fixed-rank tensor manifolds have, to our knowledge, not previously been equipped with a Riemannian metric with known geodesics, other than in the rank $1$ case [Swijsen22b, Jacobsson24].
In Riemannian optimization, line searches and gradient descents are often performed along geodesics, or along approximate geodesics, called retractions. As already mentioned, on any homogeneous manifold there is a distinguished Riemannian metric called the canonical metric. This metric has two big advantages: geodesics on a quotient can be described as geodesics on its numerator; and those geodesics are always complete, meaning they can be extended to arbitrary length. Similarly, a retraction on the quotient can be induced by a retraction on the numerator. Riemannian optimization using the canonical metric has been explored, for example, on the Grassmann [Helmke07, Sato14, Boumal15, Bendokat24], Stiefel [Edelman98, Li20, Gao22], and symplectic Stiefel [Gao21, Bendokat21] manifolds.
@@
Section˜2 briefly recaps the concepts from three different fields of study—homogeneous spaces, multilinear algebra, and algebraic geometry—that we use to construct our homogeneous structure. Sections˜3, 4, and 5 apply these to sets of tensors with fixed CP rank, multilinear rank, and TT rank respectively. The constructions are similar to each other, and the basic results are repeated without proof in sections˜4 and 5 after having been introduced in detail in section˜3.
Acknowledgements
I am grateful to my PhD examination committee, who read and gave feedback on parts of this text that were in my thesis. I am also grateful to Nick Vannieuwehoven (also committee member) and Tim Seynnaeve for fruitful discussions about algebraic geometry.
@@ 2.3 Multilinear algebra
The CP rank (often just rank) of a tensor $T\in\mathbb{R}^{n_{1}}\otimes\dots\otimes\mathbb{R}^{n_{d}}$ is the smallest $r$ such that TT is a sum of $r$ rank $1$ tensors:
@@
Figure 2: Penrose diagram for the CP decomposition. Edges are labeled with the vector space they represent, and $V_{i}=\begin{bmatrix}v_{i}^{1}&\dots&v_{i}^{r}\end{bmatrix}$. $I$ is the diagonal tensor $I_{i_{1}\dots i_{d}}=1$ if $i_{1}=\dots=i_{d}$ else $I_{i_{1}\dots i_{d}}=0$.
@@
The TT rank of a tensor $T\in\mathbb{R}^{n_{1}}\otimes\dots\otimes\mathbb{R}^{n_{d}}$ is the tuple $(s_{1},\dots,s_{d-1})$ with
@@
If $T\in\mathbb{R}^{n_{1}}\otimes\dots\otimes\mathbb{R}^{n_{d}}$ has TT rank $(s_{1},\dots,s_{d-1})$ then $T$ can be written as
@@
Such an expression is called a TT decomposition. It is illustrated in figure˜4.
Figure 4: Penrose diagram for the TT decomposition ˜2.10. Edges are labeled with the vector space they represent.
Note also that a converse of proposition˜5 is true, that a TT decomposition with inner dimensions $s_{1}$, …, $s_{d-1}$ has TT rank at most $(s_{1},\dots,s_{d-1})$.
@@ 3 The CP manifold
Note that $v_{i}^{1}$, …, $v_{i}^{r}$ are linearly independent in $\mathbb{R}^{n_{i}}$ implies $r\leq n_{i}$ for all $i$.
@@ 3.1 Smooth manifold
We now consider the subgroup of $G$ that fixes the TT defined in ˜3.3. Applying (Kruskal’s theorem)., we have the following result.
@@ 3.2 Representatives
and note that $h=(h_{1},\dots,h_{d})\in H$. Here, we see that $g_{11}$ needs to be invertible. If $g_{11}$ is not invertible, there is a permutation matrix $P$ such that $g_{i}^{\prime}=Pg_{i}$ has an invertible block $g^{\prime}_{11}$. In this way we only need to specify $g_{11}$, $g_{21}$, and $P$ for each $i$.
@@ 3.3 Riemannian manifold
For any $g\in G$, the coset $gH$ is a submanifold of $G$. Its tangent space at $g$ is called the vertical space, and is denoted $\mathcal{V}_{g}$. The orthogonal complement to the vertical space is called horizontal space, and is denoted $\mathcal{H}_{g}$. Thus $\mathfrak{h}$ and $\mathfrak{m}$ are the vertical and horizontal spaces at $1$.
@@
We now want to derive a more explicit description of the horizontal space. First, note that $\mathcal{V}_{g}=g\mathfrak{h}$, so that the vertical space consists of elements $Y=(Y_{1},\dots,Y_{d})$ where the $Y_{i}=g_{i}\begin{bmatrix}Y_{11}&Y_{12}\\ &Y_{22}\end{bmatrix}$. From now on, we suppress the $i$ superscript, but note that $Y_{11}=Y^{i}_{11}$ and the other blocks still depends on $i$. Second, if we choose a block lower triangular representative as in section˜3.2, we have
@@
But $BA$ is an $r\times r$ matrix, allowing 1/(1+BA)-1/(1+BA) to be evaluated efficiently. This trick is also useful later for evaluating the matrix exponential. For $\Gamma_{12}$, we have the following expression,
@@
has the same dot product with the $k$th column of $\begin{bmatrix}g_{11}\\ g_{12}\end{bmatrix}$ for every $i$.
@@ 3.4 Riemannian homogeneous manifold
Conversely, assume there exists an invariant subspace $\mathfrak{p}$ when $n_{i}>r$ for some $i$. Note that elements in $\mathfrak{p}$ are determined by their projection to $\mathfrak{m}$. Concretely, elements in $\mathfrak{p}$ are determined by their $r$ first columns. So consider an element $X=(X_{1},\dots,X_{d})\in\mathfrak{p}$ with
@@ 4.1 Smooth manifold
Let $E_{i}$ be the identity matrix $I_{n_{i}\times t_{i}}$ seen as an element of $\mathbb{R}^{n_{i}}\otimes\mathbb{R}^{t_{i}}$ and let $I$ be the identity matrix $I_{t_{1}\times t_{1}}$ seen as an element of $\mathbb{R}^{t_{1}}\otimes\dots\otimes\mathbb{R}^{t_{d}}$. Then define
@@ 5 The tensor train manifold
The TT decomposition is unique up to a change of basis in $\mathbb{R}^{s_{1}}$, …, $\mathbb{R}^{s_{d-1}}$. More precisely,
@@
Uschmajew and Vandereycken [Uschmajew13, proposition 3] shows that this holds for Hierarchical Tucker decompositions, of which the TT decomposition is a special case.
@@ 5.1 Smooth manifold
Let $\operatorname{ttrank}^{-1}(s_{1},\dots,s_{d-1})$ denote the set of tensors with TT rank $(s_{1},\dots,s_{d-1})$. Like $\operatorname{rank}^{-1}(r)$, its closure is an algebraic variety.
@@
From ˜2.9, it is clear that TT rank is preserved under $G$. The action is thus well-defined.
@@
The set of tensors with TT rank $(s_{1},\dots,s_{d-1})$ and multilinear rank $(s_{1},$ $s_{1}s_{2},\dots,s_{d-1})$ is a smooth homogeneous manifold,
@@ Appendix B Appendix: Counting the operations
  • no extra operations to build $A$,
@@
    • $2n_{i}r^{2}$ operations to multiply with $A$ from the left,
@@
    • $2n_{i}r^{2}$ operations to multiply with $A$ from the left.